NOAA SATOPS Morning Report: Friday, June 10, 2011


Customers or users who want to call to ask about problems or report outages can use the following contacts:

- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 161

GOES
GOES-12 experienced a cycle slip and the 2127Z image was lost. 2144Z
image was also lost during cycle slip recovery.

All other GOES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
ESPC MetOp-A Antarctica Data Acquisition (ADA) transition was
successfully executed in the early morning of June 10, 2011 (DOY 161).

All other POES operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
DOY 161 June 10, 2011 0055Z: A bad AMR calibration file prevented the
OGDR product for rev 13892 from being correctly generated. JPL provided
a corrected AMR calibration file and the OGDR product for rev 13893 was
generated later than usual. The problem has been corrected and all
subsequent OGDRs have been generated on time.

All other JASON-2 operations were nominal over the past 24 hours.
No anomalies affecting product delivery or s/c health & safety

ESPC Reported data outage and delays
DOY 160 June 9, 2011 1452Z - 1606Z: Equipment Outage: The DCS feed from
Wallops CDA to the NSOF was interrupted. Root cause under investigation.

DOY 160 June 9, 2011: GOES-13 East 1815Z image and the GOES-11 West
1730z, 1745z, 1800z images were lost when the NPORT GINI ingestors hung.
Under investigation.

DOY 160 June 9, 2011 1800Z - DOY 161 June 10, 2011 0330Z: ESPC did not
receive Meteosat-9 data. Server reboot corrected issue.

*_Data Delayed:_*

DOY 161 June 11, 2011: Windsat data from Monterey not being received
since 0930z. Monterey is investigating.

Questions or comments to Ron.Mahmot@noaa.gov

Please follow SpaceRef on Twitter and Like us on Facebook.